OpenStudy (anonymous):

Using complete sentences, explain how to graph x + 3y = 3 using the x- and y-intercepts method.

OpenStudy (mayankdevnani):

first took the value of =0

OpenStudy (mayankdevnani):

0+3y=3 3y=3

OpenStudy (mayankdevnani):

can you solve it? @mikala1

OpenStudy (anonymous):

no i dont know any thing about this tis is my fist lesson on this

OpenStudy (mayankdevnani):

y=3/3=1

OpenStudy (mayankdevnani):

so first order pair is (0,1)

OpenStudy (mayankdevnani):

ok

OpenStudy (anonymous):

and how did you come up with that could you explain the stepss please

OpenStudy (mayankdevnani):

first take the value of x=0 0+3y=3 3y=3 y=3/3 y=1

OpenStudy (mayankdevnani):

ok got it

OpenStudy (anonymous):

yes thank you

OpenStudy (mayankdevnani):

welcome